Exercise 2-13 Departmental Predetermined Overhead Rates [LO2-1, LO2-2, LO2-4]

White Company has two departments, Cutting and Finishing. The company uses a job-order costing system and computes a predetermined overhead rate in each department. The Cutting Department bases its rate on machine-hours, and the Finishing Department bases its rate on direct labor-hours. At the beginning of the year, the company made the following estimates:

Department Cutting Finishing Direct labor-hours 6,800 76,000 Machine-hours 58,400 1,600 Total fixed manufacturing overhead cost $ 380,000 $ 403,000 Variable manufacturing overhead per machine-hour $ 4.00 — Variable manufacturing overhead per direct labor-hour — $ 2.75

Required:

1. Compute the predetermined overhead rate for each department. 2. The job cost sheet for Job 203, which was started and completed during the year, showed the following:

Department Cutting Finishing Direct labor-hours 3 12 Machine-hours 85 3 Direct materials $ 780 $ 370 Direct labor cost $ 69 $ 276 Using the predetermined overhead rates that you computed in requirement (1), compute the total manufacturing cost assigned to Job 203. 3. Would you expect substantially different amounts of overhead cost to be assigned to some jobs if the company used a plantwide predetermined overhead rate based on direct labor-hours, rather than using departmental rates?
Answer #1

1) Cutting department overhead rate = 380000/58400+4 = 10.51 per machine hour

Finishing department overhead rate = 403000/76000+2.75 = 8.05 per labor hour

2) Calculate job Cost

Cutting Finishing Total
Direct material 780 370 1150
Direct labor 69 276 345
Overhead 10.51*85 = 893.35 8.05*12 = 96.6 989.95
Total 1742 743 2485

3) Yes, would be different